OpenStudy (anonymous):

Below are the steps to solve an equation: Step 1: |x − 2| + 3 = 7 Step 2: |x − 2| = 7 − 3 Step 3: |x − 2| = 4 Which of the following is a correct next step to solve the equation? x + 2 = −4 −x − 2 = 4 x + 2 = 4 x − 2 = −4 I thought it was B, but I got it wrong. Could some one please help me?

OpenStudy (ahsome):

If I said that \(|x-y|\)=z, what are the two posibilities?

OpenStudy (anonymous):

|x − 2| = 4 Remove absolute x-2=-4

OpenStudy (ahsome):

@Lyralei, there is more than 1 posibility for that

OpenStudy (anonymous):

hears all the steps Let's solve your equation step-by-step. |x−2|+3=7 Step 1: Add -3 to both sides. |x−2|+3+−3=7+−3 |x−2|=4 Step 2: Solve Absolute Value. |x−2|=4 We know either x−2=4 or x−2=−4 x−2=4(Possibility 1) x−2+2=4+2(Add 2 to both sides) x=6 x−2=−4(Possibility 2) x−2+2=−4+2(Add 2 to both sides) x=−2 Answer: x=6 or x=−2

OpenStudy (anonymous):

Okay, well that dosent work. Because I already know that x=-2 @Ahsome

OpenStudy (anonymous):

like @Ahsome said |x+y| = z have two answers x+y =z or x+y =-z

OpenStudy (ahsome):

Yes @Anastashia. What are the two posibilities?